OpenStudy (abmon98):

given that logax^2y=p and that loga(x/y^2)=q find logax and logay in terms of p and q and hence express logaxy in terms of p and q

OpenStudy (abmon98):

\[\log_{a} x^2y=p\] \[\log_{a} (x/y^2)=q\] \[\log_{a} x^2y=\log_{a} b^p\] \[x^2y=a^p\] \[\log_{a} (x/y^2)=\log_{a} a^q\] \[x/y^2=a^q\]

OpenStudy (abmon98):

@ganeshie8 i dont know how the answers are \[\log_{a} x=1/5(2p+q)\] \[\log_{a} y=1/5(p-2q)\] \[\log_{a} xy=3/5p-1/5q\] so do you have any idea on figuring out this question

ganeshie8 (ganeshie8):

we're given these : \(\large \log_a \left( x^2y\right) = p\) \(\large \log_a \left(\frac{x}{y^2}\right) = q\)

OpenStudy (abmon98):

yes

ganeshie8 (ganeshie8):

\(\large \log_a \left( x^2y\right) = p \\ \large \implies \log_a x^2 + \log_a y = p \\ \large \implies 2\log_a x + \log_a y = p ~~~~\color{red}{(1)}\)

ganeshie8 (ganeshie8):

\(\large \log_a \left( \frac{x}{y^2}\right) = q \\ \large \implies \log_a x - \log_a y^2 = q \\ \large \implies \log_a x - 2\log_a y = q ~~~~\color{red}{(2)}\)

ganeshie8 (ganeshie8):

solve \(\large \color{red}{(1)}\) and \(\large \color{red}{(2)}\)

OpenStudy (abmon98):

why in the 2nd equation is it logay^2

OpenStudy (abmon98):

is this a rule to log both sides of the fraction

ganeshie8 (ganeshie8):

yes here are the rules : \(\large \log_a \left(M*N\right) = \log_a M + \log_a N\) \(\large \log_a \left(\frac{M}{N}\right) = \log_a M - \log_a N\)

OpenStudy (abmon98):

oh i forgot that rule thank you so much \[5\log_{a} x=2p+q\] \[2p+q/5=\log_{a} x\] \[2*(2p+q/5)+\log_{a} y=p\] \[\log_{a} y=-2q+p/5\] \[\log_{a}xy= \log_{a}x+\log_{a} y\] \[(2p+q/5)-(p-2q)/5\]